ELECTROCHEMISTRY HELP????

How much electricity is required to produce a ton of lithium metal in a commercial electrolysis plant? Assume that a ton is 2000 lbs, and that the process is 100% efficient and enter the answer in Coulombs. (One pound = 0.454 kg)
Answer #1

Sol .

As 1 ton = 2000 pounds (lbs)

1 pound (lb) = 0.454 Kg

1 Kg = 1000 g

So , Mass of Lithium ( Li ) = 1 ton = 2000 lbs

= 2000 × 0.454 Kg

= 2000 × 0.454 × 1000 g

= 908000 g

As Molar Mass of Li = 6.941 g/mol

So , Moles of Li = n = Mass of Li / Molar Mass of Li  

= 908000 / 6.941 mol  

As Half cell reaction at cathode : Li+(aq) + e- ---> Li(s)

So , Electrochemical equivalent = Z = 1  

Also , Faraday's constant = F = 96500 C / mol

Therefore , Electricity Required = n × F × Z  

= ( 908000 × 96500 × 1 ) / 6.941

= 12623829419.392 C